# SpikCommander: A High-performance Spiking Transformer with Multi-view Learning for Efficient Speech Command Recognition

### v2

@@ Appendix A Appendix for SpikCommander
We apply SpecAugment (park2019specaugment) to the Mel spectrogram of input audio, incorporating both frequency and time masking to enhance robustness. For the spike trains in SHD and SSC datasets, we adopt an event drop augmentation strategy (wang2024efficient; gu2021eventdrop), which randomly removes events through two operations: drop-by-time and drop-by-neuron, each applied at a fixed ratio to introduce temporal and spatial sparsity. The effect of this augmentation is illustrated in Fig. 5 and Fig. 6, which compare examples from the GSC and SSC datasets before and after augmentation, respectively. The detailed parameter configurations for both SpecAugment and event drop are provided in Table 5.
@@
Our work is implemented using the PyTorch-based SpikingJelly (fang2023spikingjelly) framework. Additionally, we provide detailed experimental settings for the three datasets, as shown in Table 7.
@@
Datasets
Multi-view
Learning
Spiking Self-attention Mechanisms
SSA (zhou2023spikformer) SDSA (yao2023spikedriven) STASA (ours)
@@
Five Seeds Analysis. Most previous studies (wang2024efficient; wang2025efficient; liu2024lmuformer; sun2024delayed; sun2025towards; Zhou2024DirectTH) in spiking and non-spiking auditory tasks report results based on a single random seed. To ensure the reliability of our evaluation, we additionally conducted experiments with five random seeds, as summarized in Table 7. All models were trained with 100 time steps. As shown in Table 9, our results demonstrate that while smaller datasets such as SHD exhibit slightly higher variance, the overall performance remains state-of-the-art. For larger datasets like SSC and GSC, SpikCommander achieves remarkably stable accuracy with minimal standard deviation, confirming the robustness of SpikCommander. For readability and consistency, the following experiments and analyses in this paper continue to follow the single-seed configurations reported in Table 7 (corresponding to Table 1 in the main text).
@@
The results demonstrate that performance consistently improves with increasing $T$, indicating strong long-term temporal modeling capacity. On SHD, accuracy rises steadily from 83.68% (T=10T{=}10) to a peak of 96.41% (T=100T{=}100), beyond which it saturates or slightly degrades. This may be attributed to the relatively small scale of SHD, where most informative spiking cues are already captured within low resolution. For instance, SpikCommander achieves 92.39% and 93.60% accuracy at only T=40T{=}40 and T=50T{=}50, respectively, indicating that further increasing the sequence length may introduce redundant context or lead to overfitting. This observation aligns with recent works (wang2024efficient; hammouamrilearning2024; baronig2025advancing) on SHD, where competitive results are often achieved with extremely compact models, further illustrating the limited scale and relatively low complexity of the dataset.
This long-term learning ability becomes more evident on large-scale datasets such as SSC and GSC, where model performance consistently improves with increasing time steps. Notably, on SSC, performance continues to rise with longer sequences: at T=100T{=}100, the 1-block variant achieves 83.26% accuracy. When the time step increases to T=200T{=}200, its performance improves to 84.89%, achieving a relative gain of 1.63%. Further extending the sequence to T=250T{=}250 yields a final accuracy of 85.50%, with an additional gain of 0.61% over the T=200T{=}200 result. These trends indicate that SpikCommander can effectively leverage extended temporal information when available, exhibiting a strong capacity for long-term temporal modeling. On GSC, which uses Mel spectrograms as input, strong performance is already observed at short durations. For example, SpikCommander achieves 95.82% with a single block at T=40T{=}40. The performance already surpasses recent SOTA SNN models (wang2024efficient; hammouamrilearning2024; deckers2024co).) As $T$ increases, performance continues to improve steadily. At T=200T{=}200, the single-block model achieves 96.89%. When extended to T=250T{=}250, the accuracy slightly fluctuates to 96.76% (1L), indicating a saturation trend and confirming the model’s capability to stably capture long-term dependency without overfitting.
@@
The effectiveness of STASA and the multi-view design. As shown in Table 8, our proposed STASA consistently achieves the highest accuracy across all three datasets under the 1-block configuration, validating its effectiveness as a novel spiking self-attention mechanism. In particular, the introduction of multi-view learning further boosts performance for all three attention variants (SSA (zhou2023spikformer), SDSA (yao2023spikedriven), and STASA), demonstrating its general applicability and complementary benefits. Notably, STASA achieves best results in the majority of scenarios, both with and without multi-view learning. While STASA’s performance is marginally close to SSA in some cases, STASA distinguishes itself with a linear complexity of $\mathcal{O}(ND)$, offering significant computational advantages over SSA’s quadratic complexity $\mathcal{O}(N^{2}D)$. This efficiency is crucial for scaling to larger datasets and models. Moreover, STASA significantly outperforms SDSA, which also with linear complexity, further highlighting STASA’s architectural advantages. These results demonstrate STASA’s favorable trade-off between computational efficiency and representational capacity, making it a strong candidate for scalable, energy-efficient spiking self-attention mechanism for resource-constrained neuromorphic computing applications.
@@
Spiking Transformer baselines. We follow the official implementation of TIM (shen2024tim) to ensure a reliable and reproducible comparison between Spiking Transformers. Specifically, we validated our reproduction pipeline by evaluating TIM and Spikformer under the same 10 time steps setting on the SHD dataset as reported in the TIM. As shown in Table 12, our reproduced results demonstrate 85.68% for TIM (vs. 85.1% reported) and 86.54% for Spikformer (vs. 86.3% reported), which match the original, confirming the correctness of our setup. To explore model scalability across different temporal resolutions, we extended the evaluation to 40 and 100 time steps on both SHD and SSC datasets. While TIM demonstrates strong performance under shorter sequences, we observed that Spikformer and SDT maintain more consistent improvements as the sequence length increases. Therefore, in this work, we adopt Spikformer and SDT as the main comparison baselines for three benchmarks (see Table 1), while still following the implementation of TIM to ensure consistency and fairness.
@@
where $FLOP^{l}$ denotes the number of MAC operations in layer $l$, and $E_{MAC}=4.6pJ$ represents the energy cost per MAC operation on 45nm hardware (zhang2024spike).
For SNNs, the theoretical energy consumption of an SNN is usually calculated through multiplication between the number of MAC/AC operations and the energy consumption of each operation on predefined hardware (panda2020toward; zhou2023spikformer; yao2024spikedriven; zhang2024sglformer). The number of synaptic operations (SOPs) is calculated as:
@@
where $fr^{l-1}$ is the firing rate of spiking neuron layer l1l-1. $FLOP^{l}$ refers to the number of floating-point MAC operations (FLOPs) of layer $l$, and $SOP^{l}$ is the number of spike-based AC operations (SOPs). Assuming the MAC and AC operations are performed on the 45nm hardware (horowitz20141), i.e. $E_{MAC}=4.6pJ$ and $E_{AC}=0.9pJ$, the energy consumption of SpikCommander can be calculated as follows:
@@
$SOP_{Conv}$ represents the SOPs of a convolution or linear layer, and $SOP_{SSA}$ represents the SOPs of an SSA module, $FLOP_{Conv}^{1}$ represents the FLOPs of the first layer before encoding input frames into spikes. $N$ is the total number of convolution layers and linear layers, and $M$ is the number of SSA modules. During model inference, several cascaded linear operation layers such as convolution, linear, and BN layers, can be fused into one single linear operation layer (zhou2024direct), still enjoying the AC-type operations with a spike-form input tensor.
@@
We utilize the SpikingJelly framework (fang2023spikingjelly) to train our model, available at https://github.com/fangwei123456/spikingjelly.
@@
  • SpikeSCR (wang2024efficient): https://github.com/JackieWang9811/SpikeSCR
  • TIM (shen2024tim): https://github.com/BrainCog-X/Brain-Cog/tree/main/examples/TIM
  • Spikformer (zhou2023spikformer): https://github.com/ZK-Zhou/spikformer/issues/30
  • SDT (yao2023spikedriven): https://github.com/BICLab/Spike-Driven-Transformer
Our energy consumption calculation framework is based on syops-counter (chen2023training), with code available at https://github.com/iCGY96/syops-counter.
@@
SNNs are particularly appealing for their compatibility with neuromorphic hardware due to their sparse and event-driven nature. In line with recent influential SNN studies (zhou2023spikformer; yao2023spikedriven; zhou2024qkformer; xing2024spikelm) in other realms (e.g., CV, NLP), we have reported the theoretical energy consumption of SpikCommander and compared it against other SNN-based approaches. This analysis further validates the energy efficiency of our model. To bridge algorithm and hardware co-design, we are exporting our model into a Loihi-compatible .hdf5 format using Intel’s Lava framework and are also actively pursuing access to Intel’s Loihi-2 hardware (orchard2021efficient; davies2021advancing) for future deployment and empirical validation. This line of work will help validate the real-world deployment potential of our model on dedicated low-power chips.

### v3

@@ Appendix A Appendix for SpikCommander
We apply SpecAugment (Park et al. 2019) to the Mel spectrogram of input audio, incorporating both frequency and time masking to enhance robustness. For the spike trains in SHD and SSC datasets, we adopt an event drop augmentation strategy (Wang et al. 2024a; Gu et al. 2021), which randomly removes events through two operations: drop-by-time and drop-by-neuron, each applied at a fixed ratio to introduce temporal and spatial sparsity. The effect of this augmentation is illustrated in Fig. 5 and Fig. 6, which compare examples from the GSC and SSC datasets before and after augmentation, respectively. The detailed parameter configurations for both SpecAugment and event drop are provided in Table 5.
@@
Our work is implemented using the PyTorch-based SpikingJelly (Fang et al. 2023) framework. Additionally, we provide detailed experimental settings for the three datasets, as shown in Table 7.
@@
Datasets Multi-view Learning Spiking Self-attention Mechanisms
SSA (Zhou et al. 2023b) SDSA (Yao et al. 2023) STASA (ours)
@@
Five Seeds Analysis. Most previous studies (Wang et al. 2024a, 2025; Liu et al. 2024; Sun et al. 2024, 2025; Zhou et al. 2024b) in spiking and non-spiking auditory tasks report results based on a single random seed. To ensure the reliability of our evaluation, we additionally conducted experiments with five random seeds, as summarized in Table 7. All models were trained with 100 time steps. As shown in Table 9, our results demonstrate that while smaller datasets such as SHD exhibit slightly higher variance, the overall performance remains state-of-the-art. For larger datasets like SSC and GSC, SpikCommander achieves remarkably stable accuracy with minimal standard deviation, confirming the robustness of SpikCommander. For readability and consistency, the following experiments and analyses in this paper continue to follow the single-seed configurations reported in Table 7 (corresponding to Table 1 in the main text).
@@
The results demonstrate that performance consistently improves with increasing $T$, indicating strong long-term temporal modeling capacity. On SHD, accuracy rises steadily from 83.68% (T=10T{=}10) to a peak of 96.41% (T=100T{=}100), beyond which it saturates or slightly degrades. This may be attributed to the relatively small scale of SHD, where most informative spiking cues are already captured within low resolution. For instance, SpikCommander achieves 92.39% and 93.60% accuracy at only T=40T{=}40 and T=50T{=}50, respectively, indicating that further increasing the sequence length may introduce redundant context or lead to overfitting. This observation aligns with recent works (Wang et al. 2024a; Hammouamri et al. 2024; Baronig et al. 2025) on SHD, where competitive results are often achieved with extremely compact models, further illustrating the limited scale and relatively low complexity of the dataset.
This long-term learning ability becomes more evident on large-scale datasets such as SSC and GSC, where model performance consistently improves with increasing time steps. Notably, on SSC, performance continues to rise with longer sequences: at T=100T{=}100, the 1-block variant achieves 83.26% accuracy. When the time step increases to T=200T{=}200, its performance improves to 84.89%, achieving a relative gain of 1.63%. Further extending the sequence to T=250T{=}250 yields a final accuracy of 85.50%, with an additional gain of 0.61% over the T=200T{=}200 result. These trends indicate that SpikCommander can effectively leverage extended temporal information when available, exhibiting a strong capacity for long-term temporal modeling. On GSC, which uses Mel spectrograms as input, strong performance is already observed at short durations. For example, SpikCommander achieves 95.82% with a single block at T=40T{=}40. The performance already surpasses recent SOTA SNN models (Wang et al. 2024a; Hammouamri et al. 2024; Deckers et al. 2024).) As $T$ increases, performance continues to improve steadily. At T=200T{=}200, the single-block model achieves 96.89%. When extended to T=250T{=}250, the accuracy slightly fluctuates to 96.76% (1L), indicating a saturation trend and confirming the model’s capability to stably capture long-term dependency without overfitting.
@@
The effectiveness of STASA and the multi-view design. As shown in Table 8, our proposed STASA consistently achieves the highest accuracy across all three datasets under the 1-block configuration, validating its effectiveness as a novel spiking self-attention mechanism. In particular, the introduction of multi-view learning further boosts performance for all three attention variants (SSA (Zhou et al. 2023b), SDSA (Yao et al. 2023), and STASA), demonstrating its general applicability and complementary benefits. Notably, STASA achieves best results in the majority of scenarios, both with and without multi-view learning. While STASA’s performance is marginally close to SSA in some cases, STASA distinguishes itself with a linear complexity of $\mathcal{O}(ND)$, offering significant computational advantages over SSA’s quadratic complexity $\mathcal{O}(N^{2}D)$. This efficiency is crucial for scaling to larger datasets and models. Moreover, STASA significantly outperforms SDSA, which also with linear complexity, further highlighting STASA’s architectural advantages. These results demonstrate STASA’s favorable trade-off between computational efficiency and representational capacity, making it a strong candidate for scalable, energy-efficient spiking self-attention mechanism for resource-constrained neuromorphic computing applications.
@@
Spiking Transformer baselines. We follow the official implementation of TIM (Shen et al. 2024) to ensure a reliable and reproducible comparison between Spiking Transformers. Specifically, we validated our reproduction pipeline by evaluating TIM and Spikformer under the same 10 time steps setting on the SHD dataset as reported in the TIM. As shown in Table 12, our reproduced results demonstrate 85.68% for TIM (vs. 85.1% reported) and 86.54% for Spikformer (vs. 86.3% reported), which match the original, confirming the correctness of our setup. To explore model scalability across different temporal resolutions, we extended the evaluation to 40 and 100 time steps on both SHD and SSC datasets. While TIM demonstrates strong performance under shorter sequences, we observed that Spikformer and SDT maintain more consistent improvements as the sequence length increases. Therefore, in this work, we adopt Spikformer and SDT as the main comparison baselines for three benchmarks (see Table 1), while still following the implementation of TIM to ensure consistency and fairness.
@@
where $FLOP^{l}$ denotes the number of MAC operations in layer $l$, and $E_{MAC}=4.6pJ$ represents the energy cost per MAC operation on 45nm hardware (Zhang et al. 2024a).
For SNNs, the theoretical energy consumption of an SNN is usually calculated through multiplication between the number of MAC/AC operations and the energy consumption of each operation on predefined hardware (Panda et al. 2020; Zhou et al. 2023b; Yao et al. 2024; Zhang et al. 2024b). The number of synaptic operations (SOPs) is calculated as:
@@
where $fr^{l-1}$ is the firing rate of spiking neuron layer l1l-1. $FLOP^{l}$ refers to the number of floating-point MAC operations (FLOPs) of layer $l$, and $SOP^{l}$ is the number of spike-based AC operations (SOPs). Assuming the MAC and AC operations are performed on the 45nm hardware (Horowitz 2014), i.e. $E_{MAC}=4.6pJ$ and $E_{AC}=0.9pJ$, the energy consumption of SpikCommander can be calculated as follows:
@@
$SOP_{Conv}$ represents the SOPs of a convolution or linear layer, and $SOP_{SSA}$ represents the SOPs of an SSA module, $FLOP_{Conv}^{1}$ represents the FLOPs of the first layer before encoding input frames into spikes. $N$ is the total number of convolution layers and linear layers, and $M$ is the number of SSA modules. During model inference, several cascaded linear operation layers such as convolution, linear, and BN layers, can be fused into one single linear operation layer (Zhou et al. 2024a), still enjoying the AC-type operations with a spike-form input tensor.
@@
We utilize the SpikingJelly framework (Fang et al. 2023) to train our model, available at https://github.com/fangwei123456/spikingjelly.
@@
  • SpikeSCR (Wang et al. 2024a): https://github.com/JackieWang9811/SpikeSCR
  • TIM (Shen et al. 2024): https://github.com/BrainCog-X/Brain-Cog/tree/main/examples/TIM
  • Spikformer (Zhou et al. 2023b): https://github.com/ZK-Zhou/spikformer/issues/30
  • SDT (Yao et al. 2023): https://github.com/BICLab/Spike-Driven-Transformer
Our energy consumption calculation framework is based on syops-counter (Chen et al. 2023), with code available at https://github.com/iCGY96/syops-counter.
@@
SNNs are particularly appealing for their compatibility with neuromorphic hardware due to their sparse and event-driven nature. In line with recent influential SNN studies (Zhou et al. 2023b; Yao et al. 2023; Zhou et al. 2024c; Xing et al. 2024b) in other realms (e.g., CV, NLP), we have reported the theoretical energy consumption of SpikCommander and compared it against other SNN-based approaches. This analysis further validates the energy efficiency of our model. To bridge algorithm and hardware co-design, we are exporting our model into a Loihi-compatible .hdf5 format using Intel’s Lava framework and are also actively pursuing access to Intel’s Loihi-2 hardware (Orchard et al. 2021; Davies et al. 2021) for future deployment and empirical validation. This line of work will help validate the real-world deployment potential of our model on dedicated low-power chips.